OpenStudy (anonymous):

f(x) = 1/x + 5/x - 3. Find domain of the function algebraically.

OpenStudy (lucaz):

5/(x-3) ?

OpenStudy (misty1212):

set each denominator equal to zero and solve for \(x\)

OpenStudy (anonymous):

Im not sure if is \[\Large f(x) = \frac{ 1 }{ x }+ \frac{ 5 }{ x }-3\] or \[\Large f(x) = \frac{ 1 }{ x }+ \frac{ 5 }{ x-3 }\] in both cases, you have to see for what values of x the function will be undefined? is easily to see that with x=0 the function becomes undefined, and with x=3 is the same, so the domain of the function f(x) is \[\mathbb{R} - \left \{0, 3\right \} \]